If 96.5 moles of an ideal gas occupies 41.5 liters at 301 K, what is the pressure of the gas?
1
Expert's answer
2012-11-28T09:21:01-0500
PV = nRT
P = nRT/V
P = 96.5*8.31*301/41.5 = 5816,3 kPa